OpenStudy (anonymous):

a wave is modeled with the function y=(1/2)sin 3x. describe the graph of this function including its period, amplitude, and points of intersection with the x axis. I think the amp. is 1/2 and period is 2pi/3??? don't know how to find points of intersection.

OpenStudy (kropot72):

You are correct in your values for the amplitude and period. The points of intersection are found where y = 0. What are the solutions of\[0=\frac{1}{2}\sin 3x\]

OpenStudy (kropot72):

@berrypicking Are you there?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

don't know how to get soln

OpenStudy (kropot72):

\[\frac{1}{2}\sin 3x=0\ when\ 3x=0,\ \pi,\ 2\pi,\ 3\pi,\ 4\pi,\ ...........\ ,\ n\pi\] Now you need to solve for x to find each of the points of intersection with the x axis.

OpenStudy (kropot72):

Hint: The first point of intersection is when x = 0/3 = 0

OpenStudy (anonymous):

(0,0)

OpenStudy (anonymous):

( pi/3,0), (2pi/3,0) (pi,0), (4 pi/3,0) is this it

OpenStudy (kropot72):

You are correct. Good work! There is an infinite number of solutions so you should indicate this in your answer.

OpenStudy (anonymous):

thank you . I just didn't know what to make = to what.

OpenStudy (kropot72):

You're welcome :)